# InfNeRF: Towards Infinite Scale NeRF Rendering with O(log n) Space Complexity

### v1

@@ 3.2 Anti-Aliasing Rendering
As illustrated in Fig. 3(a), when the camera is far away from the scene, all spheres are relatively large, and therefore, their densities and colors are estimated by the root node, resulting in a smooth anti-aliasing effect. Conversely, as shown in Fig. 3(b), zooming in with the camera reduces the size of the spheres, which shifts the estimation to the leaf nodes, yielding a sharp and detailed image. As InfNeRF necessitates only a minimum subset of the nodes for rendering a frame, it requires loading merely number of parameters from the whole model into memory. The upper bound is reached when the camera looks at the distant horizon, as illustrated in Fig. 3(c). In this scenario, multiple-level nodes are engaged simultaneously, from the root node representing the blurring horizon to the leaf node representing foreground details.
3.3 Tree Pruning
@@
While we can build a perfect octree as in Section 3.1, this may not be efficient, particularly for non-uniformly sampled scenes. To optimize the structure, it is important to identify the essential nodes necessary for rendering all input images, effectively pruning the superfluous parts of the tree.
@@
In this work, we utilize sparse points obtained from the Structure from Motion (SfM) technique to approximate scene geometry. Each sparse point observed in images is treated as spheres, each with a radius as defined in Section 3.2. We find the corresponding nodes for all the spheres in an ideal, infinitely deep octree as described in Section 3.2. We retain only the nodes that correspond to sparse spheres and their respective parent nodes, while the others are pruned away.
@@
Compared to MegaNeRF or BlockNeRF, which split the scene uniformly based on space occupancy, our approach offers a more intelligent resource allocation strategy. It adaptively creates deeper branches in areas with rich details and shallower branches in coarse regions. In addition, our solution provides users with the flexibility to control the depth of the tree, enabling a tailored balance between reconstruction quality and computational efficiency.
3.3.1 Pruned Tree Rendering.
@@
The pruned tree structure introduces new challenges in InfNeRF rendering, as the optimal node determined by Eq. 4 may no longer exist after pruning. In such scenarios, we instead process the sampling sphere with the nearest available ancestor of the intended node.
@@
Specifically, given a sampling sphere , the rendering process recursively travels down the tree. If the sphere successfully arrives at the node of the expected level, the density and color of the sphere will be estimated by that node. However, if this process is interrupted by a pruned node, the rendering of the sphere reverts to its parent node. We outline this rendering strategy with the pseudo-code in Algorithm 1. Finally, each sampling sphere is processed by the NeRF of the located node, ensuring accurate and efficient rendering despite the pruned tree structure.
3.4 Training
To train the proposed InfNeRF, we accumulate the density and color of all spheres sampled on a ray using volume rendering [25] and compare the resulting RGB values with the pixels of the observed images using L2 loss .
@@ 3.4 Training
To remove the cloudy effect from the untrained areas, particularly in the sky, and enhance visual quality in the bird’s-eye view, we introduce a transparency loss , which uniformly samples points within the AABB and uses L1 loss to encourage their density to approach 0.
@@
To further improve the result, we also incorporate the distortion loss from [3] and a regularization loss to promote consistency across different scales by encouraging the same locations from different scales have similar density. The total loss of our training is summarized as:
3.4.1 Pyramid Supervision.
The photographs in most datasets are captured by drones from relatively close proximity, resulting a lack of multi-resolution information. While this offers sharp and detailed pixels that can well supervise the lower-level nodes, it leads to insufficient training for the higher-level nodes close to the root, which require low-frequency supervision signals.
To mitigate this issue, we augment the training set using an image pyramid constructed by repeatedly downsampling the high-resolution images. As illustatrated at Fig. 4, this pyramid supervision ensures that the upper nodes receive adequate training, effectively addressing the undertraining problem and facilitating a more balanced learning across all levels of the model.
Note that the pyramid supervision is not applicable to conventional NeRF models. The capability of InfNeRF to decompose the scene along the scale dimension is what enables it to effectively utilize multi-resolution training data. The benefit it extends beyond simple pyramid supervision. For instance, when a particular location is captured in photographs taken from varying distances, conventional NeRF models tend to compromise the low-frequency and high-frequency supervision from these images, leading to sub-optimal results. In contrast, InfNeRF naturally separates the blurry and sharp information to supervise different nodes, resulting in more accurate rendering. This concept applies vice versa as well. For instance, within a single image, detailed foreground pixels may supervise to lower-level nodes, while blurry background pixels may automatically supervise to higher-level nodes.
3.4.2 Distributed Training.
@@
Given a specific tree level , we split the octree into two parts: the upper tree which includes all nodes from level 0 to level , and the lower part which is a forest including a maximum of subtrees. During training, the upper tree is shared across all devices, and the parameters are updated concurrently. Meanwhile, individual subtrees or clusters of adjacent subtrees in the lower part, are assigned among different devices and updated independently. In other words, each device only maintains the shared upper tree and a subset of the lower subtrees, effectively reducing the VRAM footprint and communication overhead. During forward pass, if a sampling sphere descends to a subtree that is not owned by the current device, it will be forwarded by its parent node in the shared tree as demonstrated in Fig. 6. Consequently, during training, the only data transfer is the all-reduce operation on the gradient of the shared upper tree, which leads to a more efficient and scalable training process.
@@
It is worth mentioning that the training data can also be divided among devices for further acceleration. Intuitively, if a ray on a particular device does not intersect with the AABB of any subtree on that device, it would be not sensible to sample the corresponding pixel of the ray. Therefore, we find a 2D AABB for each device by projecting the sparse points inside the subtrees’ 3D AABBs back onto images. Then, during the training phase, only the pixels that fall within the 2D AABB are sampled and utilized.
4 Experiments
We show the main results in this section and present more analysis and evaluations in the supplementary material. The source code and trained models will be released to the public.
@@ 4 Experiments
Implementation Details. Inf-NeRF is implemented using the NeRFStudio framework [33]. Each node’s NeRF is an Instant-NGP implemented by tiny-cuda-nn, with default parameters such as a grid size of 2048 and 16 levels of hash tables, each with a size of . The tree is limited to 4 levels, resulting a reasonable resolution of 5cm for a scene spanning 1 km. More implementation details are provided in supplementary materials.
4.1 Experiment on Space Complexity
The crucial metric of our study is the required model size during rendering. It determines how much information must be retrieved through the network and stored in the GPU. We compare three methods with similar total model sizes, including (1) Nerfacto [33], an improved Instant-NGP [27] and whose capacity is scaled up 64 times to match our model. (2) Nerfacto-s, a 4 levels Nerfacto, representing the scale partition methods like PyNeRF [37]. (3) InfNeRF. (4) InfNeRF leaf, using the same space partition as InfNeRF, but only having leaf nodes. It can be considered as an improved version of xy grid partition.
In this experiment, a zoom-out trajectory from a close-up view to a bird’s-eye view in VGA resolution is rendered by these 4 methods, and the model size in Megabytes required for 6 keyframes is presented in Fig 6. InfNeRF exhibits a memory footprint that is only 17% of InfNeRF leaf and 12.3% of Nerfacto-s. The trajectory and the rendered images are shown in Fig. 1 and in the video of supplement material.
| Scene | Residential | Sci-Art | ||||||
| Metric | PSNR0 | PSNR | SSIM | LPIPS | PSNR0 | PSNR | SSIM | LPIPS |
| Nerfacto | 24.32 | 24.31 | 0.772 | 0.216 | 26.75 | 26.48 | 0.827 | 0.182 |
| Mega-NeRF | 22.52 | 22.85 | 0.748 | 0.265 | 24.37 | 24.22 | 0.808 | 0.202 |
| InfNeRF compact | 23.84 | 24.87 | 0.785 | 0.198 | 25.83 | 25.81 | 0.805 | 0.212 |
| InfNeRF | 24.66 | 25.46 | 0.820 | 0.158 | 27.03 | 26.96 | 0.839 | 0.1918 |
| Scene | Rubble | Building | ||||||
| Metric | PSNR0 | PSNR | SSIM | LPIPS | PSNR0 | PSNR | SSIM | LPIPS |
| Nerfacto | 24.99 | 26.09 | 0.692 | 0.336 | 22.53 | 22.94 | 0.663 | 0.325 |
| Mega-NeRF | 24.17 | 26.02 | 0.73 | 0.294 | 21.70 | 22.43 | 0.696 | 0.282 |
| InfNeRF compact | 24.62 | 27.27 | 0.731 | 0.274 | 22.52 | 24.02 | 0.703 | 0.280 |
| InfNeRF | 25.12 | 27.92 | 0.781 | 0.203 | 23.11 | 24.70 | 0.740 | 0.242 |
4.2 Experiment on Rendering Quality
Baseline. We compare the rendering quality of our approach with (1) Nerfaco [33]. (2) Mega-NeRF [36] with 25 partitions. In addition to these two methods, we also train and evaluate a smaller version of InfNeRF, called InfNeRF compact (InfNeRF-C), whose model size is similar to Mega-NeRF.
Metrics. Quantitative results are reported based on metrics such as PSNR, SSIM, and LPIPS using the Alex implementation. As zoom-out is a crucial use case in our paper, we not only report metrics at the finest resolution,denoted as PSNR0, but more importantly, the mean metrics over 6 scaling levels, denoted as PSNR, SSIM, LPIPS.
@@ 4.2 Experiment on Rendering Quality
Comparison. The PSNR0 (full resolution) and all 3 metrics for multi-resolution are listed in Table 1. The mean of 2 PSNR across the four datasets are also plotted in Fig.8. Comparing between two methods with similar rendering model sizes, InfNeRF outperforms Mega-NeRF by 2.4dB in multi-resolution PSNR. Even the InfNeRF compact with a much smaller rendering model size, still outperforms Mega-NeRF by 1.6 dB. This illustrates the overall performance of our proposed method. When comparing two models with large model sizes, InfNeRF and Nerfacto, theoretically, they should exhibit similar PSNR0. Our experiments confirm this hypothesis. InfNeRF slightly outperforms Nerfacto by 0.3 dB, demonstrating that the octree partitioning of InfNeRF does not compromise rendering quality. However, when considering multi-resolution, thanks to the LoD anti-aliasing effect, the difference enlarges to 1.3 dB. The same conclusion can be drawn from the quality comparison in Fig.9. In the last two rows of each dataset, which display the low-resolution images, we can observe that without the anti-aliasing effect, Nerfacto and Mega-NeRF generate noisy results, especially the heavy moire pattern in the left bottom dataset green region. And InfNeRF generates satisfying results at both fine and coarse levels.
It would be insightful to delve deeper into the performance of different scales. The PNSR of 6 scale of residential dataset is plotted in Fig.8, where the horizontal axis represents the scale of the image. A scale of 0 corresponds to full resolution, while a scale of 5 corresponds to resolution. In our experiment, Nerfacto and Mega NeRF reconstruct the scene with only one level, resulting in a lack of anti-aliasing effects and consequently lower PSNR at higher scale levels (lower resolution). As the resolution increased, these 2 methods’ performance improved. However, as the resolution continued to rise, the visual information surpassed the network capacity limit, leading to a drop in PSNR. Combining the aliasing and under-fit effects, the curves of the two methods form a peak at the intermediate scale. On the other hand, the two versions of InfNeRF, compact and full, distribute the model capacity to each scale intelligently, resulting in more consistent PSNR values across all scales.
5 Conclusion and Future work

### v2

@@ 3.2. Anti-Aliasing Rendering
As illustrated in Fig.Ā 3(a), when the camera is far away from the scene, all spheres are relatively large, and therefore, their densities and colors are estimated by the root node, resulting in a smooth anti-aliasing effect. Conversely, as shown in Fig.Ā 3(b), zooming in with the camera reduces the size of the spheres, which shifts the estimation to the leaf nodes, yielding a sharp and detailed image. As InfNeRF necessitates only a minimum subset of the nodes for rendering a frame, it requires loading merely number of parameters from the whole model into memory. The upper bound is reached when the camera looks at the distant horizon, as illustrated in Fig.Ā 3(c). In this scenario, multiple-level nodes are engaged simultaneously, from the root node representing the blurring horizon to the leaf node representing foreground details.
3.3. Tree Pruning
@@
While we can build a perfect octree as in SectionĀ 3.1, this may not be efficient, particularly for non-uniformly sampled scenes. To optimize the structure, it is important to identify the essential nodes necessary for rendering all input images, effectively pruning the superfluous parts of the tree.
@@
In this work, we utilize sparse points obtained from the Structure from Motion (SfM) to approximate scene geometry. Each sparse point observed in images is treated as spheres, each with a radius as defined in SectionĀ 3.2. We find the corresponding nodes for all the spheres in an ideal, infinitely deep octree as described in SectionĀ 3.2. We retain only the nodes that correspond to sparse spheres and their respective parent nodes, while the others are pruned away.
@@
Compared to Mega-NeRF or Block-NeRF, which split the scene uniformly based on space occupancy, our approach offers a more intelligent resource allocation strategy. It adaptively creates deeper branches in areas with rich details and shallower branches in coarse regions.
This approach is robust as the subtree pruning is only restricted to cases where there are no sparse points in a relatively large cube. As further detailed below, even in the rare case that a subtree is pruned by mistake, like in low texture area, the affected volume can still be well reconstructed by its parent node, albeit in a coarser manner. This ensures that, while fine details may be omitted, the essential structure is preserved.
3.3.1. Pruned Tree Rendering
@@
The pruned tree structure introduces new challenges in InfNeRF rendering, as the optimal node determined by Eq.Ā 5 may no longer exist after pruning. In such scenarios, we instead process the sampling sphere with the nearest available ancestor of the intended node.
@@
Specifically, given a sampling sphere , the rendering process recursively travels down the tree. If the sphere successfully arrives at the node of the expected level, the density and color of the sphere will be estimated by that node. However, if this process is interrupted by a pruned node, the rendering of the sphere reverts to its parent node. We outline this rendering strategy with the pseudo-code in AlgorithmĀ 1. The additional computational overhead introduced by tree traversal includes times floating-point comparisons and memory reads, which can be considered negligible when compared to the NeRF evaluation.
3.4. Training
3.4.1. Training Loss
@@
To train the proposed InfNeRF, we accumulate the density and color of all spheres sampled on a ray using volume renderingĀ (Mildenhall etĀ al., 2021) and compare the resulting RGB values with the pixels of the observed images using L2 loss . To remove the cloudy effect from the unobserved areas, particularly in the sky, and enhance visual quality in the birdās-eye view, we introduce a transparency loss , which uniformly samples points within the AABB and uses L1 loss to encourage their density to approach 0.
@@
To further improve the result, we also incorporate the distortion loss fromĀ (Barron etĀ al., 2022) and a regularization loss who measures the L2 differences of the density between levels, to promote the density consistency.The total loss of our training is summarized as:
3.4.2. Pyramid Supervision.
The images in our datasets are captured by drones at relatively close proximity, resulting in a lack of multi-resolution information necessary to supervise the higher-level nodes. To mitigate this issue, following PyNeRFĀ (Turki etĀ al., 2023)ās approach, we augment the training set using an image pyramid constructed by repeatedly downsampling the high-resolution images. The lower-resolution image with a larger pixel size, create larger sampling spheres according to Eq.Ā 3. These larger sampling spheres are sampled from higher-level nodes in the octree, effectively supervising their training. To account for the pixel imbalance across different pyramid levels, we perform uniform sampling in pixel domain, where a high-resolution image receives four times the sampling rate compared to its low-resolution counterpart. This strategy facilitates more balanced training across different levels of the image pyramid.
3.4.3. Distributed Training
@@
Given a specific tree level , we split the octree into two parts: the upper tree which includes all nodes from level 0 to level , and the lower part which is a forest including a maximum of subtrees. During training, the upper tree is shared across all devices, and the parameters are updated concurrently. Meanwhile, individual subtrees are assigned among different devices and updated independently. In other words, each device only maintains the shared upper tree and a subset of the lower subtrees, effectively reducing the VRAM footprint and communication overhead. For instance, letās consider a perfect octree with four levels. When the root node is shared , the shared upper tree contains only a tiny fraction, approximately 0.17%, of all the parameters in the entire tree. During forward pass, if a sampling sphere descends to a subtree that is not owned by the current device, it will be forwarded by its parent node in the shared tree as demonstrated in Fig.Ā 4. Consequently, during training, the only data transfer is the all-reduce operation on the gradient of the shared upper tree, which leads to a more efficient and scalable training process.
Illustration of the distributed training in InfNeRF
@@
Furthermore, the training data can also be divided among devices for further acceleration. Intuitively, if a ray on a particular device does not intersect with the AABB of any subtree on that device, it would be not sensible to sample the corresponding pixel of the ray. Therefore, we find a 2D AABB for each device by projecting the sparse points inside the subtreesā 3D AABBs back onto images. Then, during the training phase, only the pixels that fall within the 2D AABB are sampled and utilized. The experimental result of distribution training is provided in supplementary materials.
4. Experiment
@@
InfNeRF is implemented using the NeRFStudio frameworkĀ (Tancik etĀ al., 2023). Each nodeās NeRF is an Instant-NGP with default parameters such as a grid size of 2048 and 16 levels of hash tables, each with a size of . Following NeRF-wĀ (Martin-Brualla etĀ al., 2021), each image is assigned an appearance embedding vector to encode its lighting condition. More implementation details such as learning rate and experiment settings are provided in supplementary materials.
4.1. Experiment On Large-Scale Dataset
4.1.1. Dataset
To evaluate the scalability of our method, we adopt a similar approach to Mega-NeRF (Turki etĀ al., 2022), by conducting our main experiment using four real-world large-scale urban scenes dataset captured by drone from Mill 19(Turki etĀ al., 2022) and UrbanScene3D (Lin etĀ al., 2022). Each dataset consists of around two thousand 4K images, amounting to 40 gigapixels and spanning 1 km. Notably, they are approximately 200 times larger than the scenes reported in other general NeRF work. Our octree is limited to 4 levels, resulting a reasonable GSD ranges from 0.4 m of the root node to 5 cm of the leaf node. The total model size is about 3 GB.
4.1.2. Result on Space Complexity
The paramount objective of our study is to reduce the required model size in rendering, which refers to the sum of all sub NeRFsā size required for rendering one frame. It is the total information that needs to be retrieved from the clouds and stored in the local device during real-time rendering. To demonstrate the impact of our octree on the required model size, we compare InfNeRF with three other methods with similar total model sizes but with different architectures. These three methods are: (1) NerfactoĀ (Tancik etĀ al., 2023), (2) Nerfacto-s, a 4 levels Nerfacto, representing the scale partition methods like PyNeRF (Turki etĀ al., 2023), (3) InfNeRF leaf, an InfNeRF without upper tree structure, representing the space partition methods like Mega-NeRFĀ (Turki etĀ al., 2022).
In this experiment, we examine the rendering of a zoom-out trajectory, starting from a close-up view and transitioning to a birdās-eye view in VGA resolution. The trajectory and the corresponding rendered images are shown in Fig. 1. We present the required model size in megabytes for six keyframes in Fig.Ā 5. When rendering the close-up view, the space partition method InfNeRF leaf proves advantageous as it only requires a single node of the scene and has low space requirements. In contrast, the scale partition methods Nerfacto-s necessitate the detailed version of the entire scene, resulting in a higher space complexity. As the camera moves away, the space partition method InfNeRF leaf now requires all blocks to render the complete scene, while the scale partition method reduces its requirement by only rendering the low-resolution node. However, InfNeRF consistently exhibits its low space complexity in both scenarios, with a memory footprint of only 17% compared to InfNeRF leaf and 12.3% compared to Nerfacto-s.
4.1.3. Result on Rendering Quality
As zoom-out is a crucial use case in our paper, we not only report metrics at the finest resolution, denoted as PSNR0, but more importantly, the mean metrics over 6 resolution levels, denoted as PSNR, SSIM, LPIPS. We compare our approach with (1) NerfactoĀ (Tancik etĀ al., 2023), whose model size is scaled up to InfNeRF, (2) PyNeRFĀ (Turki etĀ al., 2023), also being scaled up to InfNeRF, (3) Mega-NeRFĀ (Turki etĀ al., 2022) with 25 partitions, (4) InfNeRF compact (InfNeRF-c), whose model size is similar to Mega-NeRF.
Scene | Residential | Sci-Art | Rubble | Building | ||||||||||||
Metric | PSNR0 | PSNR | SSIM | LPIPS | PSNR0 | PSNR | SSIM | LPIPS | PSNR0 | PSNR | SSIM | LPIPS | PSNR0 | PSNR | SSIM | LPIPS |
Nerfacto | 24.32 | 24.31 | 0.772 | 0.216 | 26.75 | 26.48 | 0.827 | 0.182 | 24.99 | 26.09 | 0.692 | 0.336 | 22.53 | 22.94 | 0.663 | 0.325 |
PyNeRF | 21.30 | 24.13 | 0.753 | 0.299 | 23.85 | 24.85 | 0.815 | 0.242 | 23.48 | 27.08 | 0.744 | 0.347 | 21.70 | 24.93 | 0.743 | 0.279 |
Mega-NeRF | 22.52 | 22.85 | 0.748 | 0.265 | 24.37 | 24.22 | 0.808 | 0.202 | 24.17 | 26.02 | 0.73 | 0.294 | 21.70 | 22.43 | 0.696 | 0.282 |
InfNeRF-c | 23.84 | 24.87 | 0.785 | 0.198 | 25.83 | 25.81 | 0.805 | 0.212 | 24.62 | 27.27 | 0.731 | 0.274 | 22.52 | 24.02 | 0.703 | 0.280 |
InfNeRF | 24.66 | 25.46 | 0.820 | 0.158 | 27.03 | 26.96 | 0.839 | 0.192 | 25.12 | 27.92 | 0.781 | 0.203 | 23.11 | 24.70 | 0.740 | 0.242 |
@@ 4.1.3. Result on Rendering Quality
The PSNR0 (full resolution) and all 3 metrics for multi-resolution are listed in Table 1. The mean across the four datasets are also plotted in Fig. 6. When comparing two models with similar model sizes, InfNeRF and Nerfacto, theoretically, they should exhibit similar PSNR0. Our results confirm this hypothesis. As depicted along the horizontal axis of Fig. 6, InfNeRF slightly outperforms Nerfacto by 0.3 dB, demonstrating that InfNeRFās octree partitioning and the limited memory requirement do not compromise the rendering quality. Furthermore, when considering results of multi-resolution rendering, ploted along the vertical axis, InfNeRF outperforms Nerfacto by 1.31 dB, PyNeRF by 1.01 dB and even outperform Mega-NeRF by 2.4 dB. This illustrates the crucial role of the LoD octree in anti-aliasing rendering. The same conclusion can be drawn from the quality comparison in Fig. 9. In the last three rows of each dataset, which display the low-resolution images, we can observe that without the anti-aliasing effect, Nerfacto and Mega-NeRF generate noisy results, particularly evident in the heavy moire pattern in the left bottom dataset solar panel region. In contrast, InfNeRF generates satisfactory results at both fine and coarse levels, showcasing its superiority in handling multi-resolution rendering. More detailed images are provided in Fig.Ā 10.
4.2. Experiment On MipNeRF 360 Dataset
In addition to our previous large-scale experiment, we conduct another experiment with the garden scene of MipNeRF 360 Dataset. This smaller-scale scene allows us to evaluate and compare the performance of various methods. The following methods were included in our comparison: (1) Gaussian Splatting(Kerbl etĀ al., 2023). (2) Mip-Splatting(Yu etĀ al., 2024). (3) NerfactoĀ (Tancik etĀ al., 2023). (4) PyNeRFĀ (Turki etĀ al., 2023). All NeRF based methods have similar model size. More details of the experiment setting can be found in the supplementary material.
Resolution | |||||
Gaussian Splatting | 26.52 | 25.18 | 21.58 | 19.01 | 17.266 |
Mip-Splatting | 26.87 | 27.85 | 28.83 | 28.69 | 27.76 |
Nerfacto | 23.18 | 23.5 | 24.08 | 23.32 | 22.09 |
PyNeRF | 23.79 | 24.76 | 27.04 | 27.63 | 24.82 |
InfNeRF | 26.27 | 27.04 | 28.2 | 28.87 | 27.23 |
(1) w/o Trans Loss | 25.63 | 26.45 | 27.54 | 27.92 | 26.18 |
(2) w/o Perturbation | 25.90 | 26.87 | 27.86 | 28.23 | 27.20 |
(3) w/o Upper Tree | 24.71 | 25.22 | 24.95 | 24.02 | 22.69 |
(4) w/o Pyramid | 26.12 | 23.04 | 13.00 | 8.98 | 8.52 |
(5) w 3 levels | 24.78 | 25.2 | 26.2 | 27.26 | 25.96 |
As shown in Table. 2, InfNeRF exhibits competitive outcomes in high-resolution when compared to 3DGSĀ (Kerbl etĀ al., 2023) and Mip-SplattingĀ (Yu etĀ al., 2024). It surpasses PyNeRF and Nerfacto by an margin of 3 dB. As the resolution decreases, 3DGSĀ (Kerbl etĀ al., 2023) and Nerfactoās performance start to drop. Although anti-aliasing may not be the primary focus of InfNeRF, it consistently delivers commendable results without any additional computational cost.
4.3. Ablation Study
Table. 2 also includes an ablation study of our model on garden scene. (1) The transparency loss, improves slightly PSNR, and more importantly, removes the fog blocking the bird eyes view, as shown in Fig. 8. (2) Sampling radius perturbation not only smooth the transition between different LOD levels in some dataset, as depicted in Fig. 7, but also improve the overall PSNR. (3) Training without the upper tree structure, or (4) training without image pyramid, results in a significant reduction in PSNR at the low-resolution. (5) Removing one level of the tree reduces the model size by 4, and decreases the performance by 2 dB. (6) Tree pruning reduces the size of the tree to only 19 nodes, which accounted for only 26% of the full octree.
5. Conclusion and Future work